### v2

# MUTANT: A Recipe for Multilingual Tokenizer Design

Abstract
Tokenizers play a crucial role in determining the performance, training efficiency, and the inference cost of Large Language Models (LLMs). Designing effective tokenizers for multilingual LLMs is particularly challenging due to diverse scripts and rich morphological variation. While subword methods like Byte Pair Encoding (BPE) are widely adopted, their effectiveness in multilingual settings remains underexplored. We present MUTANT, a recipe for building multilingual tokenizers, with careful vocabulary and training data design, language-aware pre-tokenization, and subword and multiword aware training. We also introduce MUTANT-Indic, a tokenizer for India-specific multilingual LLMs, that produces linguistically coherent tokens and achieves state-of-the-art performance. Evaluated across English, $22$ Indian languages and code data, our tokenizer improves the average fertility score by $39.5\%$ over LLaMA4 and by $18\%$ over Sutra (the current best). This translates to $44\%$ improvement in inference throughput over LLaMA4 while maintaining comparable performance on English and Indic benchmarks. We present detailed ablations across tokenizer training data size, vocabulary size, merging techniques, and pre-tokenization strategies, demonstrating the robustness of our design choices.
MUTANT: A Recipe for Multilingual Tokenizer Design
Souvik Rana, Arul Menezes, Ashish Kulkarni, Chandra Khatri, Shubham Agarwal Krutrim AI, Bangalore, India {souvik.rana, ashish.kulkarni, shubham.agarwal1}@olakrutrim.com
1 Introduction
Large Language Models (LLMs) (Touvron and others, 2023; Grattafiori et al., 2024; Abdin et al., 2025; Guo et al., 2025; Yang et al., 2025; Gemma Team, 2025) rely on the crucial step of tokenization, the process of converting raw text into discrete units called tokens. Among the many proposed approaches, subword tokenization schemes such as BPE (Sennrich et al., 2016a), Unigram (Kudo, 2018), WordPiece (Song et al., 2021), and their byte-level extensions have become the de facto choice. However, tokenization remains an understudied topic within the LLM literature (Dagan et al., 2024; Mielke et al., 2021), especially in multilingual settings (Petrov et al., 2023), where, skewed fertility scores across languages, often lead to concerns around fairness, high inference latency, cost and context size. For instance, with the 22 languages listed in the Eighth Schedule of the Constitution of India [^1] https://en.wikipedia.org/wiki/Languages_with_official_recognition_in_India, these issues are especially pronounced for Indic languages comprising multiple scripts and a rich morphology. A key metric for evaluating tokenizers is the “fertility score” (or token-to-word ratio) (Ali et al., 2024) where, a lower fertility score is desirable due to more efficient (and hence cheaper) LLM training and inference. Our analysis suggests that tokenizers of popular multilingual LLMs (Gemma Team, 2025; OpenAI, 2025; Meta, 2025), largely designed for English, could produce fertility scores as high as 10.5 (LLaMA-4 tokenizer for Oriya; Table 3) for Indic languages, far worse than the near-ideal scores achieved for English. This leads to longer token sequences, higher compute overheads, and poor alignment with linguistic units.
Figure 1: MUTANT: Multilingual tokenizer training recipe. Raw text is filtered and consolidated before a two-stage training process consisting of subword and multi-word learning with ablations to identify efficient design choices.
Designing an efficient tokenizer involves making careful choices around the the tokenization algorithm, size of the vocabulary (of tokens) as well as tokenizer training data. In this work, we address five core questions and investigate how to design effective multilingual tokenizers by i) examining trade-offs between low- and high-resource languages, ii) joint versus language-specific training, iii) multilingual data balancing, iv) the role of pre-tokenization, and v) the benefits of incorporating multi-word expressions via curriculum training.
Supported by controlled experiments and systematic ablations, we present MUltilingual Tokenizer optimization ANd Training (MUTANT) as a practical and reproducible recipe (see Figure 1) for building equitable, efficient, and linguistically grounded multilingual tokenizers for LLMs. We apply our recipe to train MUTANT-Indic, a state-of-the-art tokenizer for Indic LLMs, demonstrating the benefits of our methodology in a real-world multilingual setting. MUTANT-Indic combines linguistically grounded pre-tokenization with a two-stage subword–multi-word learning process (Liu et al., 2025b), yielding a compact and semantically faithful vocabulary. Figure 3 illustrates some examples where our approach avoids fragmenting common words or idiomatic phrases into unnatural subunits across different languages. We make the following contributions:
  • We propose MUTANT, a principled and general recipe for training optimal multilingual tokenizers. We systematically study the effect of vocabulary size, data distribution, language-specific pre-tokenization and multi-stage training on multilingual performance.
  • We demonstrate the effectiveness by training MUTANT-Indic, that achieves state-of-the-art (SOTA) performance for Indic languages.
  • To the best of our knowledge, we are the first to carry out a comprehensive benchmarking of tokenizers across multiple intrinsic and downstream LLM performance measures in both pretraining from scratch as well as continual pretraining settings. We will publicly release the evaluation framework to enable reproducibility and community benchmarking.
2 Related Work
Multilingual Tokenizers.
Multilingual tokenization faces challenges from script diversity, morphology, and structural variation. Comparative studies show that vocabulary size and construction strategies strongly affect performance for morphologically rich languages (Karthika et al., 2025a), while inefficiencies in underrepresented ones, such as Ukrainian, translate to higher fertility and computational costs (Maksymenko and Turuta, 2025). Tokenization also influences how multilingual models encode morphology, as demonstrated in mT5 vs. ByT5 (Dang et al., 2024). For Indic languages, tailored resources (Kakwani and others, 2020) and IndicBERT (AI4Bharat, 2022) highlight the value of domain-specific tokenization. Recent benchmarks further reveal economic implications, with BLOOM’s tokenizer achieving the best cost efficiency among popular multilingual LLMs (ADA Sci, 2024). Together, these studies show that current multilingual tokenizers fragment low-resource and morphologically rich languages, motivating approaches that combine normalization, language-tailored pre-tokenization, and multi-word learning to achieve better efficiency and fairness.
0250002500050000500007500075000100000100000125000125000150000150000175000175000200000200000Vocabulary Sizearbndevaengukamlorpasattate
Figure 2: Vocabulary size distribution across language scripts. See Appendix A for script details.
Tokenization Algorithms.
Tokenization strategies differ in both theory and practice. While alternate sub-word tokenization algorithms have been explored in the past such as WordPiece (Song et al., 2021), Unigram LM (Kudo and Richardson, 2018), Byte Pair Encoding (BPE) remains the most widely adopted. Originally developed for compression (Gage, 1994) and later adapted for neural MT (Sennrich et al., 2016b), BPE merges frequent character pairs to balance coverage with efficiency. Its variants aim to address inefficiencies: PickyBPE (Chizhov et al., 2024) discards uninformative merges to improve vocabulary utility, while Scaffold-BPE (Lian et al., 2024) iteratively prunes low-frequency scaffold tokens to reduce imbalance and enhance downstream performance. Recent extensions like SuperBPE (Liu et al., 2025a) expand beyond word boundaries, jointly learning subwords and multi-words yielding improved compression and inference efficiency in a 2-stage curriculum. BoundlessBPE (Schmidt et al., 2024), another contemporary work, relaxes the Pre-tokenization word boundary constraint in an single stage learning step. Our work compares these two recent approaches and shows that two-stage curriculum preserves subword coverage while capturing larger semantic units in morphologically rich Indian languages.
Pre-tokenization.
Pre-tokenization plays a pivotal role in shaping token boundaries, directly influencing both compression efficiency and reasoning performance (Xue and others, 2024). SentencePiece (Kudo and Richardson, 2018) introduced a language-agnostic approach by treating input as raw streams, effective for languages without whitespace boundaries. More recent approaches like BoundlessBPE (Schmidt et al., 2024) relax pre-token constraints to improve frequency distributions, while regex-based designs continue to prove crucial for capturing script-specific structures.
3 MUTANT: Tokenizer Training Recipe
Language modeling involves estimating the probability distribution over text sequences, $P(S)$, where $S$ may represent a sentence, paragraph, or document. To achieve this, the text is first converted into a sequence of discrete tokens through a tokenization function $g(S)=X=(X_{1},X_{2},\ldots,X_{n})\in V^{n}$, where $V$ denotes the vocabulary and $n$ the sequence length. Tokenizers can be open-vocabulary, ensuring any string can be represented (e.g., byte-level), or closed-vocabulary, where unseen text maps to an out-of-vocabulary symbol (e.g., word lists) (Rae et al., 2021). In our work, we adopt an open-vocabulary approach that combines byte-pair encoding (BPE) with a UTF-8 byte fallback, following  Radford et al. (2018). In this section, we present a systematic recipe for designing multilingual tokenizers that jointly addresses data curation, pre-tokenization, and subword–multiword learning. Figure 1 provides an end-to-end overview of the pipeline, illustrating how raw multilingual text is filtered and sampled, transformed through a fixed pre-tokenization stage, and used to train a two-stage tokenizer with explicit design choices on vocabulary size and allocation, detailed below.
3.1 Training Data and Vocabulary
The composition of the training data and the vocabulary size are design choices that directly impact the quality of tokens and the tokenizer efficiency.
Training data volume: Effective tokenizer training relies on a sufficiently large carefully curated multilingual corpus; beyond moderate scale, further data scaling yields diminishing returns (Reddy et al., 2025) in quality (Section 5).
as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
Size (MB) 56 562 13 4 1 148 67 83 422 30 252 60 311 34 152 82 46 144 110 0.47 38 502 486 38
# Lines (K) 65 681 19 118 2 449 135 91 545 21 273 129 337 64 257 126 59 198 139 1 69 658 773 27
Avg W/Line 51 47 41 3 43 56 37 59 59 145 41 39 35 43 33 40 45 57 36 26 70 32 32 185
Table 1: Evaluation corpus statistics across 22 Indic languages, English, and code in Section 3.4 including size (in MB), number of lines (in K) and average words per line. We report standard ISO codes here (Appendix Section A).
Data quality and diversity: We combine diverse, high-quality sources that is representative of natural language use (Hayase et al., 2024), such as, the Web, curated multilingual datasets, Wikipedia, math, code and others. The raw data is further subjected to filtering for language/script mixing, normalization, short documents (or lines) and controlled source-wise sampling (see Appendix B.1 for MUTANT-Indic).
Vocabulary size and multilingual allocation: In multilingual settings, vocabulary allocation must be explicitly language (or script)-aware to preserve subword and multi-word granularity across both high- and low-resource languages. We determine script-wise vocabulary budgets through ablations that vary allocation across scripts, identifying distributions that balance fragmentation and coverage (Figure 2). To realize these target allocations in practice, we evaluate two strategies: i) explicit merging, where we train script-specific tokenizers and concatenates their vocabularies via rule stacking; and ii) corpus-driven alignment, where we train a single tokenizer on a multilingual corpus whose data proportions are aligned with the desired script-wise vocabulary budgets.
Explicit merging introduces distributional interference and inconsistent segmentation across scripts (Table 10). In contrast, corpus-driven alignment allows the learned vocabulary to naturally converge toward the target script-wise proportions, better mirroring corpus composition (Table 9) and achieving consistently lower fertility across scripts (Table 3), outperforming both explicit merging and public baselines.
3.2 Tokenizer Training
We adopt a two-stage curriculum following the framework of Liu et al. (2025a). This procedure forms the training strategy used in MUTANT and its Indic instantiation, MUTANT-Indic.
Regex as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
GPT-2 4.36 4.72 4.67 1.57 2.88 1.32 3.95 4.12 3.47 2.47 5.95 3.17 7.08 3.30 4.86 4.37 4.44 3.28 5.97 2.71 1.30 6.53 5.61 1.29
LLaMA-4 1.83 1.74 1.99 1.54 1.56 1.33 2.17 1.83 1.36 1.36 2.15 1.56 2.24 2.27 1.61 1.59 1.65 1.47 2.51 3.60 1.45 2.07 1.83 1.47
Table 2: Fertility scores ($\downarrow$) showing LLaMA-4 regex outperforms GPT-2 in Stage-1 tokenizer training.
Tokenizer ($\downarrow$) as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
Gemma-3 2.65 1.69 2.84 1.79 1.69 1.39 2.60 2.50 1.47 1.48 3.34 1.91 3.45 2.07 2.03 2.03 4.42 2.83 3.37 5.16 2.03 2.50 2.94 1.44
GPT-OSS 2.66 2.41 3.17 1.51 1.89 1.33 2.73 2.37 1.72 1.58 3.34 2.01 3.51 2.41 2.61 2.10 6.26 2.71 3.89 13.01 1.76 3.18 3.13 1.51
LLaMA-4 4.40 2.93 3.34 1.46 2.00 1.34 2.84 3.37 1.83 1.72 4.23 2.28 4.95 2.73 2.79 2.46 10.51 3.23 4.12 9.04 2.13 5.87 4.53 1.76
Sarvam 4.24 1.91 2.92 2.14 1.85 1.66 3.01 2.11 1.53 1.91 2.53 2.11 3.19 4.60 1.94 2.35 2.43 1.67 3.78 13.07 7.62 2.49 2.63 7.93
Sutra 2.12 2.07 3.06 2.12 1.78 1.17 2.68 2.15 1.62 1.48 2.71 2.08 3.10 2.40 2.18 2.01 2.24 1.50 3.76 2.03 2.23 2.58 2.77 1.55
MUTANT-Indic 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
Table 3: Fertility score ($\downarrow$) comparison for Indic focused and good Indic support tokenizers across languages. MUTANT-Indic performs best in 20 of 24 languages. An extended version in Table 23 (Appendix).
Tokenizer ($\uparrow$) as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
Gemma-3 6.37 10.45 5.87 2.33 6.75 4.36 5.29 6.31 9.16 7.01 6.73 6.42 7.57 6.23 8.90 8.31 3.76 4.62 6.66 2.59 3.87 9.60 6.82 5.55
GPT-oss 6.36 7.35 5.27 2.77 6.04 4.55 5.02 6.68 7.83 6.54 6.74 6.11 7.43 5.34 6.94 8.04 2.65 4.83 5.79 1.03 4.46 7.56 6.41 5.28
LLaMA-4 3.84 6.05 4.99 2.85 5.70 4.53 4.84 4.69 7.37 6.03 5.33 5.39 5.26 4.71 6.49 6.84 1.58 4.05 5.45 1.48 3.69 4.10 4.43 4.54
Sarvam-2B 3.92 9.42 5.70 1.95 6.16 3.65 4.55 7.62 8.92 5.29 9.07 5.83 8.63 2.81 9.46 7.20 7.17 7.95 6.03 1.02 1.02 9.74 8.46 1.00
Sutra 8.04 8.50 5.44 1.97 6.39 5.15 4.98 7.36 8.33 7.00 8.38 5.88 8.75 5.36 8.35 8.45 7.73 8.76 6.04 6.59 3.49 9.38 8.04 5.15
MUTANT-Indic 9.12 10.15 8.18 2.84 7.86 5.44 6.29 8.95 11.01 8.59 10.30 7.80 11.33 5.67 11.11 10.39 10.07 9.40 8.70 3.60 5.40 11.32 10.70 5.55
Table 4: Bytes-per-token score ($\uparrow$) comparison for Indic focused and good support tokenizers across languages here. MUTANT-Indic performs best in 22 of 24 languages.
Stage 1 (Subword Learning): Training begins with standard byte-pair encoding (BPE) applied after whitespace pre-tokenization. This ensures that merges occur only within word boundaries, allowing the tokenizer to learn fine-grained subword units such as roots, affixes, and common morphemes. Stage 1 continues until the vocabulary reaches a pre-defined transition point $t$ (<|V|<|V|).
Stage 2 (Multi-word Learning): After reaching $t$, training resumes without whitespace constraints, allowing BPE to merge across word boundaries. This enables the formation of multi-words, frequent multiword expressions or collocations (e.g., “in the morning” in Figure 3) improving compression and reducing token counts for common phrases.
Figure 3: MUTANT-Indic (MI) captures multi-words (e.g. “wake up", “in the morning") and avoids fragmenting Indic words (see for e.g. Bengali, Tamil).
This two-stage tokenizer training is particularly effective for scripts with complex variations where, meaningful subwords are first anchored and then composed into frequent multiword units.
3.3 Pre-tokenization
As illustrated in Figure 1, pre-tokenization is a fixed, upstream component of our tokenizer design. It segments raw text before subword learning to improve token consistency and efficiency. For MUTANT-Indic, we explore regex-based, Unicode normalization, and morphology-aware strategies. Unicode-aware regex separates punctuation, handles numeric groups, and aligns tokens with semantic units, while NFKC normalization standardizes visually identical characters to reduce orthographic sparsity (Table 25). To improve robustness across Indic scripts, we explore GPT-2 pre-tokenization rules against LLaMA-4 regex in Stage 1, where the latter yields substantially lower token fragmentation and improving token-to-word ratios by 38–40% on Indic languages (Table 2).
Script-agnostic segmentation (as in LLaMA-4 regex rules) across Indic as well as non-Indic scripts helps reduce token fragmentation, while relaxing whitespace constraints in Stage 2 enables the learning of frequent multi-word expressions (see Appendix C.7). However, unconstrained merging can destabilize generation by creating sentence-crossing tokens. Enforcing additional sentence-level boundary constraints in Stage 2 regex prevents this while preserving multi-word modeling benefits.
Additionally, morphology-aware segmentation decomposes words into roots and affixes to capture recurring morphemes. Although we explore morphology-aware segmentation, integrating it into the tokenization pipeline without incurring latency overhead is challenging (Appendix C.2).
Tokenizer ($\downarrow$) as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
Gemma-3 0.63 0.59 0.87 1.31 0.91 1.06 0.94 0.76 0.83 0.93 0.81 0.89 0.73 0.81 0.76 0.83 0.44 0.89 0.84 0.59 0.99 0.45 0.67 0.85
GPT-oss 0.63 0.83 0.95 1.03 0.96 1.00 0.96 0.71 0.94 0.95 0.79 0.90 0.72 0.89 0.94 0.85 0.60 0.85 0.94 1.43 0.83 0.56 0.71 0.88
Sutra 0.55 0.74 0.93 2.09 0.92 0.89 0.96 0.68 0.92 0.91 0.67 0.94 0.65 0.92 0.84 0.82 0.24 0.51 0.91 0.26 1.10 0.47 0.59 0.90
Sarvam 0.99 0.66 0.91 1.50 1.00 1.27 1.13 0.64 0.85 1.19 0.62 0.99 0.65 2.19 0.72 0.96 0.24 0.54 0.93 1.45 3.63 0.45 0.56 4.25
MUTANT-Indic 0.45 0.60 0.65 0.94 0.78 0.85 0.82 0.54 0.68 0.80 0.53 0.76 0.50 0.91 0.61 0.67 0.18 0.45 0.66 0.45 0.72 0.38 0.44 0.86
Table 5: NSL score ($\downarrow$) comparison for Indic focused and Good Indic support tokenizers across languages here. MUTANT-Indic performs best in 23 of 24 languages. An extended version in Table 22 (Appendix).
3.4 Evaluation Framework
We develop a modular evaluation framework supporting HuggingFace [^2] https://github.com/huggingface/tokenizers, SentencePiece [^3] https://github.com/google/sentencepiece, and TikToken [^4] https://github.com/openai/tiktoken tokenizers along with a comprehensive set of intrinsic metrics (Appendix D), including Fertility score, Normalized Sequence Length (NSL), Rényi entropy and efficiency, and Bytes Per Token (BPT). All metrics are computed at the line level and aggregated to the language level. To assess tokenizer behavior in Indic use cases, we construct an evaluation set spanning 22 Indic languages, English, and code. Table 1 reports dataset statistics: text volume, number of lines, and average words per line per language. We will release both the evaluation dataset and the framework for reproducible benchmarking and fair comparison of multilingual tokenizers.
4 Experiments and Results
In our work, we compare MUTANT-Indic, trained using our recipe, against 9 tokenizers, comprising: i) Indic-focused tokenizers: tokenizers designed primarily for Indian languages: Sutra (Tamang and Bora, 2024) and Sarvam-2B (Sarvam, 2024) (referred as Sarvam). ii) Good Indic support tokenizers: multilingual tokenizers with demonstrated capabilities for Indic languages: Gemma-3-27B-it (Gemma Team, 2025) (referred as Gemma-3), GPT-OSS (OpenAI, 2025) and LLaMA-4 (Meta, 2025). iii) General tokenizers: tokenizers of widely-used general-purpose LLMs: Qwen3-32B (Qwen, 2024) (referred as Qwen-3), LLaMA-3.2-1B (Dubey and others, 2024), Mistral-Nemo (Mistral, 2024) and DeepSeek-R1 (DeepSeek, 2025).
4.1 Intrinsic Evaluation of Tokenizers
We evaluate tokenization quality using four complementary intrinsic metrics that capture efficiency, granularity, and information utilization: (i) Fertility score (Rust et al., 2021; Scao et al., 2022), measuring subword fragmentation; (ii) Normalized Sequence Length (NSL) (Dagan et al., 2024), quantifying relative compression against a base tokenizer; (iii) Rényi entropy and efficiency (Zouhar et al., 2023), assessing information density and vocabulary utilization; and (iv) Bytes per token (Kocetkov et al., 2022), reflecting memory and storage efficiency. All metrics are reported as micro-averages per line at the language level. Formal definitions are provided in Appendix D.
As shown in Table 3, MUTANT-Indic achieves state-of-the-art fertility scores across most languages, consistently yielding the lowest token-to-word ratios among nine tokenizers with strong Indic support (see Appendix Table 23 for extended comparisons). This indicates substantially reduced fragmentation, particularly for morphologically rich scripts. We also evaluate our Stage-1 tokenizer (trained with BPE on the same data and LLaMA-4 pretokenization). As shown in Table 24 in Appendix, our Stage-1 tokenizer, already achieves state-of-the-art (SOTA) performance across most Indic languages. Bytes-per-token results (Table 4) show that MUTANT-Indic achieves higher values across languages, reflecting more information-dense tokens and improved sequence compactness. Normalized Sequence Length scores (Table 5) further confirms that MUTANT-Indic produces shorter tokenized sequences relative to baseline tokenizers, indicating superior compression efficiency. Finally, Rényi entropy and efficiency results (Table 7) demonstrate that MUTANT-Indic achieves consistently higher efficiency across languages, reflecting effective and balanced utilization of the vocabulary. Collectively, these intrinsic results establish MUTANT-Indic as a robust and efficient tokenizer across diverse scripts and languages.
Dataset LLaMA-4 MUTANT-Indic
English Benchmarks
HellaSwag 0.353 0.357
CommonsenseQA 0.206 0.204
OpenBookQA 0.216 0.218
Winogrande 0.504 0.510
GSM8K 0.016 0.018
ARC Easy 0.623 0.630
ARC Challenge 0.291 0.292
MMLU 0.252 0.249
DROP 0.048 0.036
Average 0.279 0.279
Indic Benchmarks
Indic COPA 0.544 0.556
Indic Sentiment 0.524 0.551
Indic XNLI 0.347 0.346
Indic Paraphrase 0.534 0.539
MILU (Indic Multi-turn LU) 0.261 0.258
ARC Challenge (Indic) 0.236 0.244
TriviaQA (Indic) 0.268 0.262
Average 0.388 0.394
Table 6: Extrinsic evaluation of MUTANT-Indic vs LLaMA-4 on English and Indic benchmarks.
4.2 Extrinsic Evaluation
We also evaluate the downstream model performance (see Table 6) by pretraining LLaMA-3.2 1B models using two tokenizers: i) MUTANT-Indic, our proposed tokenizer optimized for morphologically meaningful segmentation in Indic and multilingual settings, and (ii) LLaMA-4 tokenizer, chosen for comparable vocabulary size and widespread use. Both models were trained on the same dataset in iso-compute setting to ensure a fair comparison. More details in the Appendix B. We find that our tokenizer shows competitive performance across the English and Indic benchmarks. We additionally report Bits-per-Character in Appendix C.6, which remains meaningful even near random accuracy. We also trained a model using the Stage-1 tokenizer, which also attains strong downstream performance. Table 26 in Appendix shows that the Stage-1 tokenizer itself constitutes a strong and competitive baseline.
The pretraining corpus (Table 17 in Appendix) balances coverage and domain diversity. It combines web-scale sources (Nemotron CC) for general context with structured data including MegaMath (Zhou et al., 2025), StackV2 (Lozhkov et al., 2024), synthetic generations, and books. Indic-language content constitutes roughly $20\%$ of the corpus, drawn from Indic CC, Wikipedia, and Sangraha Verified (Khan et al., 2024), providing sufficient signal to evaluate cross-lingual and morphologically rich representation quality.
Gemma-3 GPT-OSS LLaMA-4 Sarvam Sutra MUTANT-Indic
Entropy $\downarrow$ 20.70 20.81 21.09 20.71 20.62 20.42
Efficiency $\uparrow$ 0.22 0.19 0.14 0.21 0.23 0.28
Table 7: Rényi’s Entropy and Efficiency across top Indic tokenizers. Higher efficiency indicates better balance between vocabulary capacity and token usage.
Model TTFT (ms) $\downarrow$ OTPT (tokens/s) $\uparrow$
LLaMA-4 19.17 $\pm$ 0.15 117.99
MUTANT-Indic 18.98 $\pm$ 0.36 169.42
Table 8: Inference latency comparison of 1B models trained with LLaMA-4 and MUTANT-Indic tokenizers.
Metric ar bn deva en gu ka ml pa ta te
Data size (MB) 106 396 2200 3590 124 644 580 307 616 617
Percentage 1.12 4.18 23.25 37.94 1.31 6.81 6.13 3.24 6.51 6.52
Vocab % 2.69 6.32 20.89 32.92 2.38 7.82 6.76 4.68 7.04 8.50
Table 9: Script-specific training data size (Total corpus size 9.4 GB) and resulting vocabulary % distribution.
4.3 Impact on latency and throughput
Next, we evaluate how tokenization impacts end-to-end model efficiency. We train two 1B-parameter models under identical conditions: one with our tokenizer and one with the LLaMA tokenizer of similar vocabulary size. We then evaluate inference efficiency over 200 samples spanning Indic languages and English, with varying input lengths. Latency [^5] https://tinyurl.com/4e7nh7c8 was measured using standard metrics, including Time-To-First-Token (TTFT), Output Throughput (OTPT), and Input Sequence Length (ISL), across 200 instances ( See Appendix C.4 for details) with 5 warm-up requests and results averaged over 10 runs. Experiments were done on 8 H100 GPUs using Triton Inference Server as backend, with a maximum generation limit of 256 new tokens. Our tokenizer yields clear efficiency gains (Table 8) which stem from improved compression: shorter token sequences encode more information per token, thereby lowering per-request computation without compromising expressivity. Overall, this demonstrates that tokenizer design directly shapes not only pretraining efficiency but also real-world deployment latency, making it a critical factor for practical model performance.
4.4 Vocabulary Allocation: Explicit vs. Corpus-Driven
We compare explicit vocabulary merging with corpus-driven joint training under script-aware budget constraints. Explicitly merging script-specific tokenizers leads to fragmented segmentation and higher fertility due to cross-script interference (Table 10). In contrast, corpus-driven joint training naturally aligns vocabulary allocation with data distribution (Table 9), achieving lower fertility and outperforming merged and public baselines (Table 3). These results indicate that joint, corpus-driven training is a more effective and scalable strategy.
Tokenizer as bn hi mai mr san te
Individual 2.05 2.13 1.21 1.35 1.75 2.49 1.40
Merged 2.32 2.14 1.55 1.57 1.73 2.79 1.95
MUTANT-Indic 1.85 1.74 1.23 1.58 1.63 2.59 1.88
Table 10: Fertility comparison ($\downarrow$) between individual script tokenizers and the merged tokenizer across selected Indic languages. Lower values are better.
Dataset w/ Original LLaMA-4 w/ MUTANT-Indic
English Benchmarks
Winogrande 0.60 0.61
GSM8K 0.05 0.05
ARC Challenge 0.40 0.39
MMLU 0.32 0.29
Average 0.34 0.34
Indic Benchmarks
Indic COPA 0.58 0.56
Indic Sentiment 0.82 0.85
Indic XNLI 0.35 0.34
Indic Paraphrase 0.57 0.53
Average 0.58 0.57
Table 11: Performance comparison on English and Indic benchmarks in Continual Pretraining (CPT) setting of LLaMA-3.2 1B model (best scores in bold).
Tokenizer as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd Avg
Single-stage (200K) 1.86 1.76 2.05 1.75 1.62 1.37 2.20 1.86 1.39 1.39 2.19 1.61 2.29 2.30 1.66 1.67 1.69 1.49 2.68 3.61 1.56 2.12 1.88 1.54 1.90
Two-stage (180K/200K) 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44 1.83
Table 12: Fertility score ($\downarrow$) for single vs two-stage curriculum training (same evaluation data as Section 3.4).
Parameter as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd Average
Data size
1G 3.02 2.32 2.71 1.62 1.64 1.33 1.97 1.62 1.50 1.43 2.16 1.85 2.83 2.62 1.72 2.13 1.68 1.50 2.46 13.02 1.43 1.92 1.82 1.91 2.42
5G 1.71 1.93 2.58 1.63 1.58 1.33 2.18 1.72 1.40 1.36 2.04 1.57 2.43 2.28 1.68 1.48 1.61 1.57 2.48 4.74 1.30 2.02 1.87 1.43 1.91
10G 1.83 1.74 1.99 1.54 1.56 1.33 2.17 1.83 1.36 1.36 2.15 1.56 2.24 2.27 1.61 1.59 1.65 1.47 2.51 3.60 1.45 2.08 1.83 1.47 1.80
25G 1.75 1.84 2.56 1.62 1.57 1.33 2.15 1.78 1.39 1.36 2.04 1.56 2.32 2.23 1.67 1.47 1.63 1.55 2.45 3.92 1.31 2.01 1.86 1.34 1.86
30G 1.76 1.84 2.32 1.62 1.57 1.33 2.13 1.78 1.39 1.36 2.03 1.57 2.31 2.24 1.67 1.47 1.63 1.54 2.45 4.02 1.31 2.00 1.87 1.35 1.86
50G 1.72 1.82 2.25 1.60 1.57 1.34 2.14 1.82 1.39 1.36 2.03 1.58 2.28 2.22 1.69 1.49 1.64 1.52 2.44 4.54 1.31 2.01 1.87 1.34 1.87
Transition point
60 1.91 1.80 2.05 1.39 1.38 1.04 2.16 1.77 1.16 1.15 2.17 1.53 2.30 2.29 1.56 1.58 1.68 1.39 2.48 3.89 1.43 2.11 1.86 1.45 1.81
75 1.91 1.79 2.05 1.41 1.38 1.04 2.16 1.77 1.16 1.15 2.16 1.53 2.30 2.28 1.56 1.58 1.68 1.39 2.47 3.91 1.43 2.10 1.86 1.45 1.81
80 1.89 1.78 2.03 1.41 1.38 1.05 2.15 1.77 1.16 1.16 2.14 1.53 2.28 2.26 1.56 1.57 1.67 1.39 2.46 3.83 1.42 2.08 1.83 1.44 1.80
85 1.87 1.77 2.01 1.43 1.39 1.06 2.13 1.76 1.17 1.16 2.13 1.53 2.26 2.25 1.56 1.56 1.66 1.39 2.46 3.78 1.42 2.07 1.82 1.44 1.80
90 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44 1.83
95 1.85 1.75 1.98 1.47 1.42 1.10 2.13 1.74 1.21 1.20 2.12 1.53 2.23 2.24 1.56 1.56 1.66 1.41 2.46 3.68 1.43 2.06 1.81 1.44 1.79
Vocab size
162K/180K 1.89 1.78 2.08 1.48 1.47 1.13 2.21 1.80 1.24 1.22 2.22 1.60 2.35 2.27 1.65 1.65 1.68 1.42 2.62 3.84 1.48 2.16 1.91 1.47 1.86
180K/200K 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.27 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44 1.82
202K/225K 1.81 1.70 1.99 1.44 1.43 1.10 2.14 1.72 1.20 1.19 2.14 1.55 2.24 2.21 1.59 1.59 1.60 1.36 2.55 3.59 1.41 2.08 1.82 1.41 1.79
225K/250K 1.78 1.67 1.95 1.42 1.42 1.09 2.11 1.69 1.19 1.17 2.10 1.53 2.20 2.17 1.57 1.57 1.57 1.34 2.52 3.45 1.38 2.04 1.77 1.38 1.75
Table 13: Ablation of tokenizer training data size (in GB), transition point (as a $\%$ of vocab size 200200K), vocab size ($t=90\%$) and its impact on fertility score ($\downarrow$). We use the same evaluation data as described in Section 3.4.
4.5 Quality Analysis: Undertrained Tokens
We analyze under-trained “Glitch” tokens in our tied-embedding LLaMA-3.2-1B models trained with both the MUTANT-Indic tokenizer and a comparable BPE tokenizer of similar vocabulary size trained on the same corpus. Both tokenizers share the first 90% of the vocabulary. The MUTANT-Indic tokenizer switches to multi-word training for the last 10% whereas the base BPE tokeniser continues standard subword training. Following Land and Bartolo (2024) to construct a reference for unused embeddings, we introduced a small set of dummy tokens into the vocabulary that have zero occurrences in the training data. Their embeddings were averaged to obtain a mean reference vector. We then retrieve the top-$K$ nearest neighbors (cosine distance), which represent potential “glitch” tokens (Geiping et al., 2024). As shown in Figure 6 (in the Appendix C.3), the MUTANT-Indic tokenizer produces far fewer such glitch tokens than the base BPE tokenizer. These results suggest that incorporating multi-words promotes more efficient utilization of the vocabulary, while purely subword-based tokenizers overfit in the long tail, yielding a higher proportion of under-trained tokens.
4.6 Can we replace Opensource model tokenizer with MUTANT-Indic?
Following ReTok (Gu et al., 2024), we replace the tokenizer of a pre-trained LLaMA-3.2-1B model (denoted LLaMA-3.2-Orig(Grattafiori et al., 2024) with MUTANT-Indic (referred as LLaMA-3.2-MUTANT-Indic). Let $V_{\text{orig}}$ and $V_{\text{MUTANT-Indic}}$ denote their corresponding vocabularies. For a token $t\in V_{\text{MUTANT-Indic}}$, we initialize its embedding $E_{\text{init}}(t)$ as: if $t\in V_{\text{orig}}\cap V_{\text{MUTANT-Indic}}$, then $E_{\text{init}}(t)=E_{\text{orig}}(t)$, its embedding from the pretrained model, otherwise, if $t\in V_{\text{MUTANT-Indic}}\setminus V_{\text{orig}}$ and decomposes under the original tokenizer into $(t_{1},\ldots,t_{k})$, then $E_{\text{init}}(t)=\tfrac{1}{k}\sum_{i=1}^{k}E_{\text{orig}}(t_{i})$.
We then continually pretrained the LLaMA-3.2-MUTANT-Indic model, keeping just the embedding and LM head layers trainable, on a 40B-token corpus comprising English, Indic, code, and mathematics (see Appendix for details). As seen in Table 11, the LLaMA-3.2-MUTANT-Indic model performs competitively with the original LLaMA-3.2-ORIG. This suggests that, in addition to pretraining-from-scratch settings, an optimized multilingual tokenizer, such as MUTANT-Indic, could also be leveraged in opensource models through CPT (Continual Pretraining (Chen et al., 2024)) leading to significant throughput gains (see Table 8).
5 Ablation studies
Two-Stage vs. One-Stage: Controlling Vocabulary
Recently, BoundlessBPE (Schmidt et al., 2024) also explored a one-stage training paradigm in which pre-tokenization is governed by a fixed regular expression, enabling the direct learning of multiword units in a single pass for English language. Our approach instead introduces a two-stage procedure. We replicate the one-stage setup of BoundlessBPE using its released regex and compare against our strategy. Our method consistently achieves lower fertility across the top 10 Indic languages and English (Table 12). Overall, the comparison highlights a clear trade-off: while one-stage methods capture surface-level patterns indiscriminately, our two-stage design balances efficiency and linguistic integrity by decoupling subword and multiword learning.
Dataset Size
Similar to Reddy et al. (2025), we study the effects of scaling training data in Stage 1 of our training. Table 13 shows that the performance plateaus after 10G of data.
Transition Point
We ablate the transition point $t$ (Section 3.2) at which training shifts from subword to cross-word merges. Varying $t$ reveals a clear trade-off: early transitions favor frequent multiword expressions but weaken morphological coverage, while late transitions preserve subwords at the cost of longer sequences. Across Indic and non-Indic languages, intermediate values of 85-90% $t$ yields the best balance, improving token efficiency and cross-lingual consistency (Table 13).
Vocabulary Size
Vocabulary size strongly influences tokenization-model efficiency with trade-offs. Smaller vocabularies yield finer subword units that generalize well to unseen words but with increased compute costs. Larger vocabularies shorten sequences by encoding frequent forms as single tokens, but waste capacity on rare items, inflate embeddings and softmax layers (Shazeer et al., 2017), and have bias toward high-resource languages, hurting multilingual balance. With the same transition point at 90%, we found no significant impact on fertility scores beyond 200K (Table 13).
Effect of Normalization
Prior works show that unicode normalization is crucial for multilingual settings (Karthika et al., 2025b), which reduces token fragmentation and inflated vocabulary size. Table 25 (in Appendix) shows that NFKC yielded marginal but consistent gains by unifying character forms. Accordingly, we adopt NFKC to reduce variability and improve tokenizer robustness.
6 Conclusion
In this work, we revisit tokenization as a central design choice for multilingual LLMs and introduce MUTANT, a principled and general recipe for multilingual tokenizer training. We demonstrate the effectiveness of this framework with MUTANT-Indic, which achieves SOTA performance across Indic languages through systematic analyses of vocabulary size, data distribution, language-specific pre-tokenization, and multi-stage learning. Extensive evaluation across intrinsic efficiency metrics, downstream tasks, ablations, and inference latency, in both pretraining-from-scratch and continual pretraining settings, highlights consistent gains in efficiency and deployment cost. We will additionally publicly release our evaluation framework to support reproducibility and future research.
Limitations
While MUTANT-Indic achieves strong intrinsic efficiency gains and practical throughput improvements, our study has several limitations. First, although we evaluate across 22 Indic languages, English, and code, the analysis is restricted to these language families, and extending the framework to other low-resource or non-Indic languages remains future work. Second, downstream evaluations are limited to models up to 1B parameters due to compute constraints; while prior scaling-law results suggest these gains extrapolate to larger models, we do not validate this empirically. Third, although we analyze morphology-aware pre-tokenization, we do not integrate it into the final tokenizer due to its high inference latency, leaving efficient implementations to future work. Finally, while our evaluation framework standardizes intrinsic metrics, it does not capture all linguistic dimensions relevant to downstream tasks.
Ethical Considerations
Ethics Statement
This work focuses on the responsible development of multilingual tokenization methods for Indian languages. We did not collect or utilize any sensitive or Personally Identifiable Information (PII). AI tools (like ChatGPT) were used to refine text writing and section phrasing in the paper. All external datasets, libraries, and tools employed in this work are appropriately acknowledged through citations. Since the study did not involve personal, medical, or otherwise sensitive information, formal IRB approval was not required. Throughout the process, we aimed to minimize biases that could disadvantage low-resource languages. We provide our exhaustive study to advance the development of inclusive and efficient multilingual language models.
Reproducibility Statement
To promote transparency and reproducibility, we will release the artifacts publicly to benchmark performance of Indian tokenizers, along with detailed documentation. Detailed records of experimental setups, hyperparameters, and evaluation protocols are maintained to allow replication of our results with the implementation details in the Appendix. In addition, we provide ablation studies to facilitate fair benchmarking and enable future research on Indian and multilingual tokenization.
Acknowledgement
We sincerely thank the Krutrim leadership for their continued support in conducting this research. We would also like to thank Aditya Kallappa, Gautam Rajeev, Guduru Manoj, Shaharukh Khan and Neel Rachamalla for their help in different parts of this work, and the Research team at Krutrim for their valuable suggestions and insightful discussions.
Appendix
Appendix A Language Details
We provide more details about the 22 languages and corresponding scripts considered in our work in Table 14. Additionally, ISO code mapping is also provided in Table 15.
Family Script Languages
Indo-Aryan Devanagari Hindi, Marathi, Maithili, Dogri, Konkani, Sanskrit, Nepali, Kashmiri
Bengali (bn) Assamese, Bengali
Gurmukhi (pa) Punjabi
Arabic (ar) Urdu, Sindhi
Dravidian Kannada (kn) Kannada
Malayalam (ml) Malayalam
Tamil (ta) Tamil
Telugu (te) Telugu
Tibeto-Burman Devanagari Bodo
Meitei Mayek Manipuri (Meitei Mayek script)
Austroasiatic Ol Chiki (sat) Santali
Table 14: Linguistic composition of the 22 scheduled Indian languages analyzed in this work, with their corresponding scripts.
Code Language Code Language Code Language
as Assamese bn Bengali brx Bodo
doi Dogri gu Gujarati hi Hindi
kn Kannada ks Kashmiri gom Konkani
mai Maithili ml Malayalam mni Manipuri
mr Marathi ne Nepali or Odia
pa Punjabi san Sanskrit sat Santali
snd Sindhi ta Tamil te Telugu
ur Urdu
Table 15: Mapping of ISO codes to corresponding 22 Indic languages.
Appendix B Implementation
B.1 Tokenizer implementation
Our training code for the tokenizer is based on the open implementation of SuperBPE [^6] https://github.com/PythonNut/superbpe/tree/main using HuggingFace library (Jain, 2022). We first curate a multilingual corpus of $\sim$50G from OLMo (OLMo et al., 2025), Wikipedia [^7] https://en.wikipedia.org/wiki/, books, PDFs, Common Crawl, and the Sangraha dataset (Khan et al., 2024) where we sample our tokenizer training data from varied sources to get enough representation of the different languages individually. English data was mostly sampled from OLMo and code data from Stackv2 (Lozhkov et al., 2024). The common crawl corpus was language-identified using fastText LID, deduplicated using MinHash, filtered for boilerplate and HTML artifacts, normalized using Unicode NFKC and script-validated to reduce cross-script contamination, similar to FineWeb (Penedo et al., 2024). The final MUTANT-Indic tokenizer uses a shared vocabulary of 200200K tokens, distributed across language scripts and is trained on 1010GB of multilingual high quality data.
In our work, we also explored merging tokenizers based on the default priority based BPE in SentencePiece [^8] https://github.com/google/sentencepiece. While we explored implementing the multi-word two stage curriculum in the SentencePiece, we found that it was not trivial. On the other hand, HuggingFace showed issues with the merging strategy. We thus relied on different implementations for different approaches.
B.2 Training details
We provide more details about our training setup as discussed in Section 4.2. Each model was trained for 50B tokens under matched hyperparameters (learning rate, batch size, training steps), aligning FLOPs to isolate tokenizer effects. The evaluation was performed using lm-eval-harness (Gao et al., 2024) across standard English benchmarks (MMLU, GSM8K, Winogrande, TriviaQA, HellaSwag, ARC, OpenBookQA, CommonsenseQA, DROP) and Indic benchmarks (IndicCOPA, IndicSentiment, IndicXParaphrase, IndicXNLI (Doddapaneni et al., 2023), ARC Challenge Indic (Sarvam AI, 2025), and MILU Verma et al. (2024)). We report EM for GSM8K and TriviaQA, F1 for DROP, and Accuracy for other benchmarks. Shot settings were fixed per task: $25$-shot for ARC/ARC Challenge Indic, $10$-shot for HellaSwag, $5$-shot for MMLU, GSM8K, and TriviaQA, and zero-shot for the remainder. This setup allows a direct assessment of how tokenizer design influences pretraining efficiency, semantic representation, and generalization across English and Indic tasks.
Tokenizer Architecture Parameters Data Size (B ) Learning Rate Train Steps Context Length Batch Size Vocab Size
LLaMA-4 LLaMA-3.2 1B 53.24 $5\times 10^{-5}$ 68000 4096 192 201134
MUTANT-Indic LLaMA-3.2 1B 53.18 $5\times 10^{-5}$ 68000 4096 192 200008
Table 16: Pretraining configuration for different tokenizers.
Category Sources Percentage (%) Token Count (B)
Web Nemotron CC 30 15
Math MegaMath 15 7.5
Code StackV2 15 7.5
Synthetic New Generations 10 5
Books Archive 10 5
Indic Indic CC 8 4
Indic Indic Wiki 4 2
Indic Sangraha Verified 8 4
Total 100 50
Table 17: Pretraining corpus distribution across domains and token count. Indic content is emphasized to reflect multilingual objectives.
Appendix C Additional Discussion
C.1 Mismatch Between Loss and Task Performance
Although toknizers, incorporating multi-word often show slightly higher loss (Liu et al., 2025a) during training compared to models using traditional atomic tokenizers like SentencePiece/BPE, this does not necessarily translate to worse downstream performance. We hypothesize that this is due to two complementary factors. First, the introduction of longer or multi-word tokens such as “to the" or “as well as" increases the number of semantically overlapping candidates, making the model’s prediction space less sharply peaked. This means the model may distribute probability across several plausible completions (e.g., “to", “to the", “to be"), thereby lowering the maximum assigned probability to the correct token and inflating the cross-entropy loss. In contrast, other BPE tokenizers often yield only one atomic candidate for such function words, allowing sharper predictions with lower loss. Second, MUTANT-Indic tokenizes text into fewer, more meaningful units, so when computing the average loss per token, each mistake contributes more heavily to the total. As a result, although the model learns more compact and generalizable representations, its token-level loss appears higher. This creates a divergence between model loss and real-world task accuracy, indicating that traditional loss curves may underrepresent the representational efficiency and practical utility of compositional tokenizers like MUTANT-Indic.
C.2 Morphologically grounded token splitting
We investigate the impact of incorporating morphological information into tokenization for Indic languages (Brahma et al., 2025). The approach involves pre-processing text with a morphology analyzer to segment words into morphemes prior to training. This experiment focuses on languages in the Devanagari script.
We compare two variants: Tokenizer A, trained on raw text, and Tokenizer B, trained on morphologically segmented text using morphology analyzer  (Kunchukuttan, 2020). At inference time, Tokenizer B requires the same pre-processing for consistency. Tokenizer B exhibits more semantically coherent multi-words, reflecting meaningful morpheme combinations (Figure 5, 5). This promotes better generalization across related forms and reduces the raw token-to-word ratio, as morpheme-based units are more compressible. Sample outputs (Figures 5, 5) illustrate the contrast between surface-level splits and linguistically aligned segmentations.
Despite these gains, we do not adopt this approach in our final tokenizer. The primary limitation is latency, as the pipeline requires both language identification and morphological analysis. For completeness, we evaluated a Hindi morphology-aware tokenizer augmented with a morphological analyzer (Kunchukuttan, 2020) combined with language identification (LID) [^9] https://pypi.org/project/langdetect/. We performed inference on approximately 4-5 MB of Hindi text and measured throughput over 10 runs (with 5 warm-up runs), comparing against our MUTANT-Indic tokenizer. Our tokenizer achieved 194K tokens/sec, whereas the morphology-aware tokenizer achieved 90K tokens/sec, representing a 53.28% reduction in throughput. This slowdown arises entirely from the additional LID and morphology-analysis stages, underscoring the efficiency advantages of our approach even when compared to linguistically informed baselines. Extending robust analyzers across all Indic languages also introduces engineering overhead and brittle dependencies. Nevertheless, morphology-aware tokenization remains a promising direction if fast, reliable analyzers become widely available.
Figure 4: Tokenized output of morph-aware tokenizer
Figure 5: Tokenized output of non morph-aware tokenizer
Figure 6: Trend of potential glitch tokens in upper 20K of vocabulary for different K.
C.3 More on Glitch tokens
For each tokenizer, we vary $K\in\{10,50,100,150,\dots,400\}$ to select the top-$K$ embeddings closest to a reference vector derived from artificially unused tokens in the vocabulary (Land and Bartolo, 2024; Geiping et al., 2024). For the MUTANT-Indic tokenizer, we count the number of multi-word tokens within the top-$K$. For the BPE variant, we count tokens with IDs >180,000>180{,}000, which corresponds to the upper 20K of the vocabulary. Both tokenizers share the first 180K IDs; the difference lies in how the final 20K IDs are utilized: MUTANT-Indic allocates this space for frequent multi-word tokens, while the BPE tokenizer continues learning subwords. This design choice allows MUTANT-Indic to more effectively utilize the tail of the vocabulary for meaningful units, whereas the BPE tokenizer exhibits overfitting in low-frequency subwords. The trend of these counts across different top-$K$ values is visualized in Figure 6. As $K$ increases, the fraction of multi-word tokens in MUTANT-Indic remains low but stable, while the BPE variant consistently shows a higher fraction of under-trained subwords, indicating overfitting in the residual vocabulary space.
C.4 More on Latency and Throughput Evaluation
To obtain reliable latency and throughput measurements, we constructed a 200-example multilingual inference set intended to approximate realistic LLM workloads. The set contains diverse sentence-completion style prompts representative of common generation patterns. We include 20 inputs per language across English and nine major Indic languages, ensuring balanced coverage of script diversity, lexical variation, and syntactic complexity. Table 18 presents the token-length distribution of these examples under both the LLaMA-4 tokenizer and our MUTANT-Indic, allowing a controlled comparison of inference efficiency across tokenization schemes.
Table 18: Token-length statistics for the 200-example inference set. We report min, p75, p90, p99, maximum, and average token lengths.
Tokenizer min p75 p90 p99 max avg
Llama-4 288 805 1157 2583 2869 784
MUTANT-Indic 178 440 541 654 676 379
C.5 Details About Baseline Tokenizers
Tokenizer fertility is shaped by multiple factors including training data distribution, vocabulary construction, and underlying algorithmic choices, yet publicly available documentation on these aspects is often limited. Table 19 summarizes the vocabulary sizes, training methodologies, and any disclosed data distributions for all baseline tokenizers considered in our study.
Tokenizer Vocab Size Training Algorithm / Framework Data Distribution
DeepSeek-R1 128K BPE (undisclosed variant) Not publicly disclosed
Gemma-3 262K SentencePiece 140+ languages
GPT-OSS 200K o200k_harmony (TikToken variant) Not publicly disclosed
LLaMA-3.2-1B 128K BPE / SentencePiece-based Not publicly disclosed
LLaMA-4 200K BPE Not fully disclosed
Mistral-Nemo 131K Tekken tokenizer (TikToken-based) 100+ languages; multilingual + code
Qwen-3 151K Byte-level BPE Not publicly disclosed
Sarvam 68K Not publicly disclosed Not publicly disclosed
Sutra 256K SentencePiece (unigram/BPE hybrid) Balanced multilingual; uniform sampling
Table 19: Summary of baseline tokenizers and publicly available training details.
C.6 Bits-per-Character Evaluation on Multiple Choice Benchmarks
In addition to task-level accuracy, we evaluate the language modeling quality of the models using bits-per-character (BPC) measured on the gold answer continuations of benchmark examples. Following Heineman et al. (2025), we compute BPC as:
$\text{BPC}(y|x)=-\frac{1}{|y|\log 2}\sum_{t=1}^{|y|}\log P(y_{t}\mid x,y_{<t}),$ (1)
where $x$ denotes the prompt, $y$ is the gold continuation (correct answer text), and $|y|$ is the number of character in the continuation. Lower values indicate better modeling of the target continuation.
This metric provides a continuous measure of modeling quality that is independent of the discrete multiple-choice selection. It allows comparison of tokenizers through the likelihood assigned to the correct answers in benchmark datasets, where accuracy is near chance.
English Benchmarks
Table 20 reports BPC values on several English multiple-choice benchmarks. Across these tasks, the model trained with the MUTANT-Indic tokenizer consistently achieves lower BPC compared to the model trained with the LLaMA-4 tokenizer, indicating improved likelihood modeling of the correct answer continuations.
Tokenizer ARC-C ARC-E OBQA HellaSwag CSQA
MUTANT-Indic 1.0284 1.2368 1.5774 0.9867 2.3445
LLaMA-4 1.2575 1.6037 1.8345 0.9983 2.7398
Table 20: Bits-per-character (lower is better) on English multiple-choice benchmarks. ARC-C: ARC-Challenge, ARC-E: ARC-Easy, OBQA: OpenBookQA, CSQA: CommonsenseQA.
Indic Benchmark
We further evaluate BPC on translated ARC-Challenge benchmarks across ten Indic languages. Results are shown in Table 21. The MUTANT-Indic tokenizer generally produces lower BPC values across the evaluated languages, suggesting improved modeling efficiency for the gold answer continuations.
Tokenizer bn gu hi kn ml mr or pa ta te
MUTANT-Indic 1.2747 1.3480 1.3066 1.2113 1.0298 1.3205 1.1975 1.2054 0.9317 1.1535
LLaMA-4 1.2756 1.4275 1.3254 1.3193 1.2264 1.4368 1.3694 1.3935 1.0508 1.2889
Table 21: Bits-per-character (lower is better) on ARC-Challenge across Indic languages.
C.7 Stage-2 Regex Boundary Constraints
Stage 1 pre-tokenization relies on established regex-based segmentation strategies commonly used in multilingual tokenizers (e.g., GPT-2 and LLaMA-style rules). These patterns separate punctuation, normalize whitespace boundaries, and produce stable token units prior to subword learning.
Stage 2 instead applies additional regex constraints that regulate subword merges during tokenizer training. The objective is to enable learning of frequent multi-word expressions while preventing unstable token formations.
Specifically, the Stage 2 regex enforces: (i) sentence boundary preservation, preventing merges across sentence delimiters; (ii) explicit handling of Indic punctuation such as the Devanagari markers ‘|” and ‘||” as hard boundaries; and (iii) controlled segmentation of numeric sequences.
Appendix D Metrics Definitions
Here, we discuss the different intrinsic metrics used in our evaluation framework.
D.1 Token-to-Word Ratio
The Token-to-word ratio measures the average number of tokens required to represent a single word. It captures the degree of segmentation induced by a tokenizer and is particularly informative for morphologically rich languages where excessive fragmentation increases sequence length. We report this metric to evaluate whether tokenizers balance compact representations with sufficient linguistic coverage.
D.2 Bytes-per-token
Bytes-per-token quantifies the average number of raw text bytes contained in a token. Since scripts differ substantially in character set size and encoding, this metric provides a language-agnostic measure of efficiency. Higher values indicate that tokens encode more information per unit, which reduces sequence length. We include this metric to enable direct comparison of tokenizers across writing systems.
D.3 Normalized Sequence Length
Normalized sequence length measures the average length of tokenized sequences relative to a chosen base tokenizer. Instead of reporting absolute sequence lengths, this metric highlights how much longer or shorter sequences become when compared to an established reference. It enables fairer cross-tokenizer comparisons since raw lengths can vary significantly across languages and corpora. A normalized value greater than one indicates that the tokenizer produces longer sequences than the baseline, while a value less than one reflects more compact tokenization. We include this metric to directly assess relative efficiency in sequence compression.
D.4 Reyni’s Efficiency
Rényi’s entropy measures the uncertainty of token distributions induced by a tokenizer, extending Shannon entropy by allowing different orders to emphasize frequent or rare tokens. A tokenizer with a very large vocabulary may contain many infrequent tokens that are poorly utilized, while a very small vocabulary forces overuse of common tokens. Entropy therefore reflects how effectively the vocabulary is allocated. To complement this, Rényi’s efficiency normalizes entropy with respect to vocabulary size, providing a scale-invariant view of how well the vocabulary capacity is utilized. Together, these metrics characterize both the distributional balance of tokens and the comparative efficiency of different vocabulary scales.
Appendix E Extended Results
In the main paper, due to space constraints, we limited the number of tokenizers presented. Here, we provide an extended list including all of our baseline tokenizers. Table 22 and 23 provides the NSL scores and fertility scores of the other tokenizers considered in our work. Table 25 shows the fertility scores for different normalization techniques in our MUTANT-Indic tokenizer. Additionally, we also compare the performance of two models trained with only Stage-1 and the 2-stage curriculum tokenizer in Table 26.
Table 22: Comparison of NSL scores (Base LLaMA-4) for different tokenizers across all languages.
Tokenizer ($\downarrow$) as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
DeepSeek-R1 0.83 0.97 1.25 1.03 1.29 0.98 1.28 1.48 1.59 1.29 1.41 1.34 1.52 0.99 1.49 1.61 0.67 1.41 1.19 0.69 1.34 0.82 1.34 1.21
Gemma-3 0.63 0.59 0.87 1.31 0.91 1.06 0.94 0.76 0.83 0.93 0.81 0.89 0.73 0.81 0.76 0.83 0.44 0.89 0.84 0.59 0.99 0.45 0.67 0.85
GPT-OSS 0.63 0.83 0.95 1.03 0.96 1.00 0.96 0.71 0.94 0.95 0.79 0.90 0.72 0.89 0.94 0.85 0.60 0.85 0.94 1.43 0.83 0.56 0.71 0.88
LLaMA-3.2-1B 1.90 2.71 1.08 1.02 1.36 0.99 1.22 2.91 1.47 1.36 3.30 1.16 3.25 1.92 1.41 1.44 1.48 2.45 1.19 1.34 1.33 2.11 3.01 1.58
LLaMA-4 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Mistral-Nemo 1.00 0.95 1.06 1.15 1.07 1.06 1.09 1.09 1.12 1.08 0.91 1.08 0.95 0.95 1.13 1.21 1.57 0.98 1.04 1.34 1.20 0.63 0.64 0.95
Qwen-3 1.68 2.37 1.78 1.11 1.85 1.03 1.72 2.59 2.65 2.16 2.69 1.97 2.57 1.72 2.35 2.47 1.19 2.37 1.92 0.96 1.37 1.63 2.45 1.63
Sutra 0.55 0.74 0.93 2.09 0.92 0.89 0.96 0.68 0.92 0.91 0.67 0.94 0.65 0.92 0.84 0.82 0.24 0.51 0.91 0.26 1.10 0.47 0.59 0.90
Sarvam 0.99 0.66 0.91 1.50 1.00 1.27 1.13 0.64 0.85 1.19 0.62 0.99 0.65 2.19 0.72 0.96 0.24 0.54 0.93 1.45 3.63 0.45 0.56 4.25
MUTANT-Indic 0.45 0.60 0.65 0.94 0.78 0.85 0.82 0.54 0.68 0.80 0.53 0.76 0.50 0.91 0.61 0.67 0.18 0.45 0.66 0.45 0.72 0.38 0.44 0.86
Table 23: Fertility scores across tokenizers and languages. Lower is better.
Tokenizer ($\downarrow$) as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
DeepSeek-R1 3.54 2.88 4.23 1.53 2.66 1.34 3.68 4.92 3.02 2.49 6.01 3.21 7.95 2.67 4.17 3.97 7.13 4.48 5.07 6.12 2.82 4.92 6.13 2.17
Gemma3 2.65 1.69 2.84 1.79 1.69 1.39 2.60 2.50 1.47 1.48 3.34 1.91 3.45 2.07 2.03 2.03 4.42 2.83 3.37 5.16 2.03 2.50 2.94 1.44
GPT-OSS 2.66 2.41 3.17 1.51 1.89 1.33 2.73 2.37 1.72 1.58 3.34 2.01 3.51 2.41 2.61 2.10 6.26 2.71 3.89 13.01 1.76 3.18 3.13 1.51
Llama-3.2-1B 8.44 8.08 3.64 1.51 2.92 1.35 3.46 9.95 2.74 2.70 14.44 2.79 16.26 5.31 3.90 3.52 15.68 7.88 4.86 12.15 2.85 12.25 13.68 2.73
LLaMA-4 4.40 2.93 3.34 1.46 2.00 1.34 2.84 3.37 1.83 1.72 4.23 2.28 4.95 2.73 2.79 2.46 10.51 3.23 4.12 9.04 2.13 5.87 4.53 1.76
Mistral-Nemo 4.28 2.82 3.52 1.75 2.12 1.41 3.08 3.63 2.05 1.82 3.84 2.48 4.82 2.67 3.10 2.97 16.92 3.04 4.34 12.16 2.51 3.67 3.71 1.65
Qwen3-32B 7.47 7.11 6.10 1.68 4.05 1.41 5.08 8.87 4.86 3.70 11.48 4.53 12.77 4.76 6.56 6.10 12.37 7.60 8.04 8.81 2.95 9.69 11.10 2.90
Sarvam-2B 4.24 1.91 2.92 2.14 1.85 1.66 3.01 2.11 1.53 1.91 2.53 2.11 3.19 4.60 1.94 2.35 2.43 1.67 3.78 13.07 7.62 2.49 2.63 7.93
Sutra 2.12 2.07 3.06 2.12 1.78 1.17 2.68 2.15 1.62 1.48 2.71 2.08 3.10 2.40 2.18 2.01 2.24 1.50 3.76 2.03 2.23 2.58 2.77 1.55
MUTANT-Indic 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
Regex as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
MUTANT-Indic (Stage-1) 1.83 1.74 1.99 1.54 1.56 1.33 2.17 1.83 1.36 1.36 2.15 1.56 2.24 2.27 1.61 1.59 1.65 1.47 2.51 3.60 1.45 2.07 1.83 1.47
MUTANT-Indic (Stage-2) 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
Table 24: Fertility scores ($\downarrow$) comparing our tokenizer after Stage 1 and Stage 2 curriculum. Stage-1 tokenizer corresponds to: cleaned corpus, standard BPE, LLaMA-4 regex, identical vocabulary and data size. The only difference between this baseline and our final MUTANT-Indic is the Stage-2 curriculum phase.
Tokenizer as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
NFC 1.8520 1.7449 2.0412 1.4658 1.4520 1.1167 2.1741 1.7664 1.2250 1.2042 2.1845 1.5761 2.3025 2.2421 1.6273 1.6241 1.6464 1.3915 2.5859 3.7170 1.4515 2.1226 1.8754 1.4371
NFD 1.8518 1.7454 2.0413 1.4665 1.4521 1.1168 2.1661 1.7667 1.2252 1.2044 2.1905 1.5765 2.3019 2.2487 1.6274 1.6246 1.6465 1.3917 2.5864 3.7170 1.4523 2.1227 1.8757 1.4377
NFKC 1.8512 1.7430 2.0409 1.4647 1.4520 1.1155 2.1738 1.7644 1.2239 1.2041 2.1812 1.5762 2.2991 2.2327 1.6258 1.6234 1.6420 1.3884 2.5855 3.7172 1.4505 2.1200 1.8724 1.4369
Table 25: Fertility scores with NFC, NFD, NFKC normalization for all languages.
Table 26: Comparison of downstream performance between MUTANT-Indic (Stage-1) and MUTANT-Indic (Stage-2).
English Benchmarks Indic Benchmarks
Dataset MUTANT-Indic-Stage-1 MUTANT-Indic-Stage-2 Dataset MUTANT-Indic-Stage-1 MUTANT-Indic-Stage-1
HellaSwag 0.348 0.357 Indic COPA 0.556 0.556
CommonsenseQA 0.193 0.204 Indic Sentiment 0.557 0.551
OpenBookQA 0.214 0.218 Indic XNLI 0.366 0.346
Winogrande 0.515 0.510 Indic Paraphrase 0.562 0.539
GSM8K 0.021 0.018 MILU (Indic Multi-turn LU) 0.265 0.258
ARC Easy 0.625 0.630 ARC Challenge (Indic) 0.247 0.244
ARC Challenge 0.279 0.292 TriviaQA (Indic) 0.268 0.262
MMLU 0.255 0.249
DROP 0.042 0.036
Average 0.277 0.279 Average 0.403 0.394